# Test-Time Adaptation with State-Space Models

### v1

@@ C.1 Yearbook and FMoW
[34, 28] Normalization statistics during test-time adaptation are a running estimates of both the training data and the incoming test statistics. No hyperparameter optimization is necessary here.
@@
[37] Like in BN, the normalization statistics are based on both training and test set. As in Wang et al. [37], we use the same optimizer settings for test-time adaptation as used for training, except for the learning rate that we find via grid search on $\{1e^{-3},1e^{-4},1e^{-5},1e^{-6},1e^{-7}\}$. For both yearbook and FMoW, Adam optimizer [21] is used.
@@
[38] We use the same optimizer as used during training (Adam optimizer [21]). For hyperparameter optimization we follow the parameter suggestions by Wang et al. [38] and conduct a grid search for the learning rate ($\{1e^{-3},1e^{-4},1e^{-5},1e^{-6},1e^{-7}\}$), EMA factor ($\{0.99,0.999,0.9999\}$) and restoration factor ($\{0,0.001,0.01,0.1\}$). We follow [38] by determining the augmentation confidence threshold as a function of the 5 % percentile for the softmax prediction confidence on the source images from the source model. Note that this requires access to the source data.
STAD-Gaussian
We initialize the mixing coefficients with $\pi_{t,k}=\frac{1}{K}\forall t,k$, the transition covariance matrix with $\bm{\Sigma}^{\mathrm{trans}}=0.01\times\mathbf{I}$ and the emission covariance matrix with $\bm{\Sigma}^{\mathrm{ems}}=0.5\times\mathbf{I}$. The prototypes at time $t=1$ are initialized with the source weights. We found a normalization of the representations to be also beneficial for STAD-Gaussian. Note that despite normalization, the two models are not equivalent. STAD-Gaussian models the correlation between different dimensions of the representations and is therefore more expressive, while STAD-vMF assumes an isotropic variance.
@@
We initialize the mixing coefficients with $\pi_{t,k}=\frac{1}{K}\forall t,k$ and the prototypes at time $t=1$ with the source weights. For yearbook, we employ class specific noise parameters initialized with $\kappa_{k}^{\mathrm{trans}}=100$ and $\kappa_{k}^{\mathrm{ems}}=100$. For FMoW, we found a more restricted transition noise model beneficial. We follow suggestions by Gopal and Yang [16] to keep noise concentration parameters fixed instead of learning them via maximum likelihood in order to maintain a regularization term. The noise parameters remain $\kappa^{\mathrm{trans}}=1000$ and $\kappa^{\mathrm{ems}}=100$ throughout adaptation.
C.2 CIFAR-10-C
For the experiments on CIFAR-10-C, we construct a gradual distribution shift setting by increasing the corruption severity sequentially from level 1 to level 5. We adapt each model separately for each of the 15 corruption types. The source model is a WideResNet-28 [42] from RobustBench [9]. CIFAR-10-C is a well studied benchmark in CTTA and thus we take the hyperparameter settings of baseline methods reported in previous work [38].
BN
As in Section C.1, normalization statistics during test-time adaptation are a running estimates of both the training statistics and the incoming test statistics. No hyperparameter optimization is required.
TENT
As in [38], we use Adam optimizer with learning rate 1e-3.
CoTTA
We follow [38] and use Adam optimizer with learning rate 1e-3. The EMA factor is set to 0.999, the restoration factor is 0.01 and the augmentation confidence threshold is 0.92.
STAD-vMF
We initialize the mixing coefficients with $\pi_{t,k}=\frac{1}{K}\forall t,k$, the transition concentration parameter with $\kappa^{\mathrm{trans}}=100$, and the emission concentration parameter with $\kappa^{\mathrm{ems}}=100$. The prototypes at time $t=1$ are initialized with the source weights.
@@ Appendix D Additional Results
D.1 STAD is more robust to class imbalance
We also investigate robustness to class imbalance through an experiment on the binary Yearbook classification task. In this experiment, each batch is sampled to maintain a fixed class share. Figure 5 presents adaptation accuracy across different class proportions. All baseline models struggle to outperform the source model when dealing with highly imbalanced classes (i.e. class proportion of 10%). However, STAD demonstrates adaptation gains significantly earlier than the baseline methods.
Figure 5: Adaptation performance for different degrees of class imbalance: Accuracy is the average performance when both the female and male classes are kept at low proportions (e.g. at 10%). STAD demonstrates greater robustness to class imbalanced test batches.
D.2 Runtime
In Table 4, we present the wall clock time per batch for the Wild-Time datasets, with figures shown relative to the source model. BN and TENT perform adaptation the fastest. STAD requires a fraction of time compared to CoTTA, which updates all model parameters. Note that CoTTA and TENT benefit from highly efficient back-propagation code, whereas our code base has not been optimized for maximum speed. Optimizing efficiency of our code could potentially make STAD even faster.
Table 4: Relative runtime per batch with respect to the source model
@@ D.2 Runtime
STAD-vMF 2.5 53.3
STAD-vMF + BN 2.5 53.3
STAD-Gaussian 3.3 -
STAD-Gaussian + BN 3.3 -

### v2

@@ C.3 Implementation Details and Hyperparameters
(Schneider et al., 2020; Nado et al., 2020) Normalization statistics during test-time adaptation are a running estimates of both the training data and the incoming test statistics. No hyperparameter optimization is necessary here.
@@
(Wang et al., 2021) Like in BN, the normalization statistics are based on both training and test set. As in Wang et al. (2021), we use the same optimizer settings for test-time adaptation as used for training, except for the learning rate that we find via grid search on $\{1e^{-3},1e^{-4},1e^{-5},1e^{-6},1e^{-7}\}$. Adam optimizer (Kingma & Ba, 2015) is used. For CIFAR-10-C, we follow the hyperparameter setup of Wang et al. (2022) and use Adam optimizer with learning rate $31e-$ 3.
@@
(Wang et al., 2022) We use the same optimizer as used during training (Adam optimizer Kingma & Ba (2015)). For hyperparameter optimization we follow the parameter suggestions by Wang et al. (2022) and conduct a grid search for the learning rate ($\{1e^{-3},1e^{-4},1e^{-5},1e^{-6},1e^{-7}\}$), EMA factor ($\{0.99,0.999,0.9999\}$) and restoration factor ($\{0,0.001,0.01,0.1\}$). Following Wang et al. (2022), we determine the augmentation confidence threshold by the 5% percentile of the softmax prediction confidence from the source model on the source images. For the well-studied CIFAR-10-C dataset, we follow the setting of Wang et al. (2022) and use Adam optimizer with learning rate $31e-$ 3. The EMA factor is set to $0.999$, the restoration factor is $0.01$ and the augmentation confidence threshold is $0.92$.
SHOT
(Liang et al., 2020) We perform a grid search for the learning rate over $\{1e-3,1e-4\}$ and for $\beta$, the scaling factor for the loss terms, over $\{0.1,0.3\}$.
SAR
(Niu et al., 2023) We conduct a grid search over the learning rate selecting among $\{1e-2,1e-3,1e-4,0.00025\}$. Like the authors, we compute the $E_{0}$ threshold as a function of number of classes $0.4\times\ln K$, use SDG, a moving average factor of 0.9, and the reset threshold of 0.2. The updated layers include all of batch, group, and layer normalization where present in the source model.
LAME
(Boudiaf et al., 2022) The only hyperparameter is the choice of affinity matrix. Like Boudiaf et al. (2022) we use a $k$-NN affinity matrix and select the number of nearest neighbours among $\{1,3,5\}$.
T3A
(Iwasawa & Matsuo, 2021) We test different values for the hyperparameter $M$. The $M$-th largest entropy values are included in the support set used for computing new prototypes. We test the values $\{1,5,20,50,100,\texttt{None}\}$. None corresponds to no threshold, i.e. all samples are part of the support set.
@@
The hyperparameters are the initialization values of the transition concentration parameter $\kappa^{trans}$, emission concentration parameter $\kappa^{ems}$ and the sliding window size $s$. We chose the concentration parameters from $\{100,1000\}$. Tab. 6 lists employed settings. We use a default window size of $s=3$. For Yearbook, we employ class specific noise parameters $\kappa_{k}^{\mathrm{trans}}$ and $\kappa_{k}^{\mathrm{ems}}$ as discussed in Sec. 3.3. For the other datasets, we found a more restricted noise model beneficial. Particularly, we use global concentration parameters, $\kappa^{\mathrm{trans}}$ and $\kappa^{\mathrm{ems}}$, and follow suggestions by Gopal & Yang (2014) to keep noise concentration parameters fixed instead of learning them via maximum likelihood (see line 17 and 18 in Algorithm 1). Keeping them fix acts as a regularization term as it controls the size of the cluster (via $\kappa^{\mathrm{ems}}$) and the movement of the prototypes (via $\kappa^{\mathrm{trans}}$). Low concentration values generally correspond to more adaptation flexibility while larger values results in a more conservative and rigid model.
STAD-Gauss
We initialize the mixing coefficients with $\pi_{t,k}=\frac{1}{K}\forall t,k$, the transition covariance matrix with $\bm{\Sigma}^{\mathrm{trans}}=0.01\times\mathbf{I}$ and the emission covariance matrix with $\bm{\Sigma}^{\mathrm{ems}}=0.5\times\mathbf{I}$. We found a normalization of the representations to be also beneficial for STAD-Gauss. Note that despite normalization, the two models are not equivalent. STAD-Gauss models the correlation between different dimensions of the representations and is therefore more expressive, while STAD-vMF assumes an isotropic variance.
Table 6: Hyperparameters employed for STAD-vMF
Dataset $\kappa^{trans}$ $\kappa^{ems}$
Yearbook (covariate shift) 100 100
Yearbook (+ label shift) 1000 100
EVIS (covariate shift) 1000 1000
EVIS (+ label shift) 1000 100
FMoW-Time (covariate shift) 100 100
FMoW-Time (+ label shift) 1000 100
CIFAR-10.1 (covariate shift) 1000 1000
CIFAR-10.1 (+ label shift) 1000 100
ImageNetV2 (covariate shift) 100 1000
ImageNetV2 (+ label shift) 1000 100
CIFAR-10-C (covariate shift) 1000 100
@@ Appendix D Additional Results
D.1 Single Sample Adaptation
Table 7: Adaptation accuracy on temporal distribution shift with single sample adaptation (batch size 1) for both covariate shift and additional online label shift: Table shows values as plotted in Fig. 5. Most methods collapse when only provided with one sample per adaptation step. STAD can improve upon the source model in 4 out of 6 scenarios.
Yearbook EVIS FMoW
Model covariate shift + label shift covariate shift + label shift covariate shift + label shift
Source 81.30 $\pm$ 4.18 56.59 $\pm$ 0.92 68.94 $\pm$ 0.20
BN 73.32 $\pm$ 6.90 73.32 $\pm$ 6.90 11.12 $\pm$ 0.97 11.12 $\pm$ 0.97 3.46 $\pm$ 0.03 3.46 $\pm$ 0.03
TENT 61.33 $\pm$ 9.42 61.45 $\pm$ 9.46 10.80 $\pm$ 0.84 10.78 $\pm$ 0.81 4.00 $\pm$ 0.87 3.99 $\pm$ 0.87
CoTTA 55.91 $\pm$ 5.26 56.71 $\pm$ 6.12 10.04 $\pm$ 0.08 9.88 $\pm$ 0.34 3.42 $\pm$ 0.14 3.42 $\pm$ 0.03
SHOT 53.71 $\pm$ 3.77 51.48 $\pm$ 1.68 20.23 $\pm$ 1.40 16.36 $\pm$ 1.67 3.84 $\pm$ 0.21 4.01 $\pm$ 0.44
SAR 73.32 $\pm$ 6.89 73.38 $\pm$ 7.01 11.12 $\pm$ 0.97 11.12 $\pm$ 0.97 3.46 $\pm$ 0.03 3.46 $\pm$ 0.03
T3A 83.51 $\pm$ 2.54 83.44 $\pm$ 2.60 57.63 $\pm$ 0.77 57.40 $\pm$ 0.76 66.78 $\pm$ 0.24 66.87 $\pm$ 0.27
LAME 81.29 $\pm$ 4.18 81.30 $\pm$ 4.18 56.59 $\pm$ 0.91 56.59 $\pm$ 0.91 68.94 $\pm$ 0.20 68.94 $\pm$ 0.20
STAD-vMF 84.32 $\pm$ 2.03 81.49 $\pm$ 4.23 56.15 $\pm$ 0.98 58.02 $\pm$ 0.77 68.88 $\pm$ 0.29 71.22 $\pm$ 0.40
D.2 STAD in Combination with BN
Table 8: We explored whether STAD, which adapts the classifier, can be effectively combined with TTA methods like BN adaptation, which targets the feature extractor. The results are mixed. On the covariate shift of Yearbook, combining the two methods improves performance beyond what each achieves individually. However, on other datasets, the combination generally results in decreased performance.
Yearbook EVIS FMoW
Model covariate shift + label shift covariate shift + label shift covariate shift + label shift
Source 81.30 $\pm$ 4.18 56.59 $\pm$ 0.92 68.94 $\pm$ 0.20
BN 84.54 $\pm$ 2.10 70.47 $\pm$ 0.33 45.72 $\pm$ 2.79 14.48 $\pm$ 1.02 67.60 $\pm$ 0.44 10.14 $\pm$ 0.04
STAD-vMF 85.50 $\pm$ 1.34 84.46 $\pm$ 1.19 56.67 $\pm$ 0.82 62.08 $\pm$ 1.11 68.87 $\pm$ 0.06 86.25 $\pm$ 1.18
STAD-vMF + BN 86.20 $\pm$ 1.23 69.96 $\pm$ 0.39 44.23 $\pm$ 2.88 15.18 $\pm$ 1.68 66.97 $\pm$ 0.46 9.26 $\pm$ 1.97
STAD-Gauss 86.22 $\pm$ 0.84 84.67 $\pm$ 1.46
STAD-Gauss + BN 86.56 $\pm$ 1.08 70.12 $\pm$0.33
D.3 Cluster Visualization
Figure 7: t-SNE visualization of the representation space of FMoW-Time (year 2013) under joint covariate and label shift: We visualize the cluster structure in representation space. Colors indicate ground truth class labels for the 10 most common classes. Adapting with BN destroys the cluster structure, resulting in inseparable clusters. In contrast, STAD operates on linearly separable representations.
D.4 Runtime
Table 9: Relative runtime per batch compared to the source model: BN and T3A are the fastest for adaptation, while STAD takes significantly less time than CoTTA, which updates all model parameters. Note that baselines like TENT and CoTTA benefit from highly optimized backpropagation code, whereas our codebase has not been optimized for speed. Improving the efficiency of our code would make STAD likely faster.
@@ D.4 Runtime
SHOT 1.3 6.3
SAR 1.5 7.1
T3A 1.1 1.8
LAME 1.2 2.9
STAD-vMF 2.5 30.8
STAD-Gauss 3.3 -
D.5 Sensitivity to Hyperparameters
Table 10: Sensitivity of concentration parameters $\kappa^{\mathrm{trans}}$ and $\kappa^{\mathrm{ems}}$ for STAD-vMF on Yearbook (uniform label distribution) in accuracy: In general, high concentration parameters model more compact clusters ($\kappa^{\mathrm{ems}}$) and limited prototype movement ($\kappa^{\mathrm{trans}}$). Low concentration parameters allow for loose clusters and faster dynamics.
Value $\kappa^{\mathrm{trans}}$ $\kappa^{\mathrm{ems}}$
50 85.5034 $\pm$ 1.3031 85.5022 $\pm$ 1.3036
100 85.5034 $\pm$ 1.3031 85.5034 $\pm$ 1.3031
500 85.5034 $\pm$ 1.3031 85.5043 $\pm$ 1.3019
1000 85.5034 $\pm$ 1.3031 85.5058 $\pm$ 1.3001
5000 85.4980 $\pm$ 1.3099 85.5058 $\pm$ 1.3001
Table 11: Sensitivity of sliding window parameter $s$ for STAD-vMF on Yearbook (uniform label distribution): A smaller $s$ reduces the influence of past prototypes, while a larger $s$ extends the considered history, giving more weight to past prototypes. However, this comes at the cost of increased computational burden, as runtime scales linearly with window size. In our experiments, we use a default window size of $s=3$. The results suggest that giving more weight to past inferred prototypes may further improve performance.
$s$ Accuracy
3 85.4975 $\pm$ 1.34
5 85.5022 $\pm$ 1.30
7 85.5029 $\pm$ 1.31